OpenStudy (anonymous):

Find the solution set for y = x + 4, given the replacement set {(-2, -2), (-1, -2), (0, 4), (1, 5), (2, 6)}.

OpenStudy (anonymous):

{(0, 4), (1, 5), (2, 6)} {(2, -2), (-1, -2), (0, 4)} {(-2, -2), (0, 4), (1, 5), (2, 6)} {(-1, -2), (0, 4), (1, 5), (2, 6)} answers

OpenStudy (ivettef365):

substitute the points on the equation to see which one is true example point (0,4) (x,y) equation: y = x + 4 4 = 0 + 4 4 = 4 true, so (0,4) is good